# SEMF: Supervised Expectation-Maximization Framework for Predicting Intervals

### v2

@@ 5.2 Missing Data
The results of models with 50% missing data are in Table 3. Note that relative metrics for SEMF are compared with the best result from any baseline imputer on that metric, regardless of how the imputer performed on the other metrics. Overall, the results are less consistent than with complete data. MultiXGBs maintained strong performance on some datasets, such as naval_propulsion_plant, but declined on others, like cpu_activity. MultiETs exhibited narrow predictions, improved absolute PICP in some cases, such as energy_efficiency, and improved relative point predictions. Conversely, MultiMLPs, while offering narrow prediction intervals, had higher PICP, indicating decreased predictive power and increased model uncertainty.
Table 3: Test results for all models with missing data at 95% quantiles aggregated over five seeds. For each metric, the mean and standard deviation of the performance across the seeds are separated by ±. Performance over the baseline is highlighted in bold.
@@
space_ga -5%±6% 4%±2% -12%±7% 0.86±0.02 2.05±0.16 0.22±0.02 -7%±5% -9%±3% 0.45±0.07
cpu_activity -5%±21% 12%±9% -37%±52% 0.89±0.07 0.76±0.24 0.14±0.05 -1%±24% -17%±14% 0.83±0.08
naval_propulsion_plant 105%±33% -16%±7% 56%±6% 0.70±0.07 0.62±0.06 0.18±0.02 -14%±22% -29%±39% 0.77±0.06
miami_housing 16%±14% -5%±6% 10%±18% 0.82±0.07 0.77±0.15 0.10±0.02 -14%±12% -18%±11% 0.72±0.11
kin8nm -5%±2% 3%±2% -14%±5% 0.88±0.01 2.46±0.07 0.48±0.01 -3%±3% -4%±5% 0.40±0.02
concrete_compressive_strength 2%±3% 8%±4% -10%±4% 0.77±0.04 1.47±0.03 0.30±0.01 -16%±14% -20%±16% 0.58±0.09
cars -15%±15% 26%±10% -63%±29% 0.85±0.04 1.58±0.18 0.27±0.03 -30%±25% -35%±27% 0.66±0.17
energy_efficiency 10%±29% 11%±6% -24%±48% 0.92±0.03 0.77±0.19 0.22±0.05 -186%±296% -153%±210% 0.91±0.09
california_housing 26%±8% -2%±3% 21%±7% 0.88±0.03 1.67±0.12 0.40±0.03 -4%±4% -5%±5% 0.69±0.05
airfoil_self_noise -2%±21% -3%±4% -4%±19% 0.73±0.05 1.59±0.33 0.34±0.07 -37%±34% -43%±40% 0.41±0.13
QSAR_fish_toxicity -2%±3% 0%±5% -6%±8% 0.73±0.03 1.70±0.13 0.26±0.02 -4%±5% -2%±4% 0.36±0.05
@@
space_ga 4%±4% -5%±1% 7%±4% 0.88±0.02 2.23±0.10 0.24±0.01 -12%±4% -14%±5% 0.40±0.06
cpu_activity -16%±12% -5%±2% -20%±21% 0.93±0.01 1.07±0.20 0.20±0.04 -30%±26% -54%±22% 0.79±0.11
naval_propulsion_plant 58%±5% -17%±5% 46%±3% 0.82±0.05 1.38±0.10 0.41±0.03 -36%±23% -93%±53% 0.71±0.06
miami_housing 20%±16% -10%±5% 21%±14% 0.87±0.05 1.15±0.26 0.14±0.03 -27%±12% -37%±12% 0.67±0.12
kin8nm 5%±3% -8%±2% 12%±4% 0.87±0.02 2.55±0.10 0.50±0.02 -11%±3% -13%±3% 0.32±0.03
concrete_compressive_strength 1%±6% -14%±4% 13%±2% 0.80±0.02 1.83±0.06 0.37±0.01 -35%±23% -50%±28% 0.49±0.07
cars -23%±7% -8%±6% -35%±16% 0.83±0.04 1.41±0.15 0.24±0.02 -33%±20% -31%±20% 0.66±0.10
energy_efficiency -15%±18% 0%±10% -36%±47% 0.85±0.06 0.51±0.05 0.15±0.01 -73%±82% -84%±72% 0.94±0.04
california_housing 21%±7% -7%±2% 22%±5% 0.91±0.02 2.19±0.13 0.52±0.03 -14%±3% -17%±4% 0.61±0.06
airfoil_self_noise -7%±20% -20%±2% 8%±18% 0.76±0.03 1.96±0.39 0.42±0.08 -65%±60% -93%±78% 0.33±0.12
QSAR_fish_toxicity 10%±8% -9%±3% 16%±4% 0.82±0.04 2.12±0.20 0.33±0.03 -8%±9% -10%±10% 0.39±0.10
@@
space_ga -11%±6% -15%±2% 1%±7% 0.68±0.04 1.44±0.09 0.16±0.01 -40%±26% -34%±17% 0.21±0.25
cpu_activity -50%±18% 22%±16% -193%±132% 0.86±0.11 0.97±0.42 0.18±0.08 -15%±21% -23%±19% 0.75±0.14
naval_propulsion_plant -52%±21% -14%±27% -102%±62% 0.63±0.20 1.17±0.28 0.34±0.08 -108%±74% -183%±218% -0.03±0.53
miami_housing -33%±17% 13%±11% -92%±78% 0.85±0.06 0.99±0.32 0.12±0.04 -25%±9% -28%±12% 0.67±0.11
kin8nm -11%±6% -5%±2% -8%±6% 0.77±0.01 1.76±0.09 0.35±0.02 -16%±9% -15%±10% 0.37±0.04
concrete_compressive_strength -15%±14% -13%±31% -10%±54% 0.60±0.25 1.11±0.59 0.22±0.12 -12%±6% -12%±7% 0.54±0.06
cars -51%±5% 43%±8% -211%±38% 0.93±0.02 1.74±0.22 0.29±0.04 -28%±14% -26%±15% 0.66±0.11
energy_efficiency -13%±20% 9%±17% -30%±28% 0.66±0.15 0.33±0.15 0.09±0.04 1%±10% -1%±18% 0.95±0.03
california_housing -17%±7% 4%±4% -28%±16% 0.88±0.03 1.84±0.17 0.44±0.04 -5%±3% -3%±4% 0.61±0.06
airfoil_self_noise 2%±18% -2%±6% -1%±23% 0.75±0.08 1.15±0.41 0.25±0.09 -15%±9% -9%±9% 0.55±0.20
QSAR_fish_toxicity -17%±5% 6%±7% -30%±12% 0.76±0.04 1.82±0.18 0.28±0.03 -13%±7% -14%±9% 0.34±0.07
@@ 5.3 Discussion
The results indicate that specific models, particularly MultiXGBs, perform strongly with complete data. However, their robustness diminishes with increasing amounts of missing data. Further, we observe that MultiETs can produce narrow prediction intervals at the expense of reduced coverage. In contrast, despite the experiment design prioritizing interval estimation over point accuracy, MultiMLPs also deliver good point predictions. Precise choices, such as processing the columns separately, using the same hyper-parameters for both the complete and incomplete data, and fixing the same number of latent nodes per input ($m_{k}$), might have further constrained the models’ ability to tailor their performance to varying dataset complexities.
The stability of SEMF across different seeds, as indicated by the absolute PICP and NMPIW, contrasts with the significant variability observed in baseline models, suggesting that SEMF offers a more reliable performance framework. The method’s sampling operation, akin to cross-validation, helps combat overfitting, ensuring robust results despite variations in the data. Nonetheless, most models do not achieve the desired 95% prediction intervals. This shortfall could be addressed by integrating conformal prediction techniques with the intervals generated by SEMF, enhancing the reliability and coverage of prediction intervals across different datasets and missing data scenarios [34].
@@ 6 Conclusion
This paper introduces the Supervised Expectation-Maximization Framework (SEMF), a novel model-agnostic approach for generating prediction intervals in datasets with missing values. SEMF draws from the EM algorithm for supervised learning to devise latent representations that produce better prediction intervals than quantile regression. Due to SEMF’s iterative simulation technique, training and inference can be done with complete and incomplete data. A comprehensive set of 330 experimental runs on 11 datasets with three different model types showed that SEMF is superior to quantile regression, particularly for models such as XGBoost, which intrinsically lacks latent representations. This research underscores the potential of SEMF in various application domains and opens new avenues for further exploration for supervised latent representation learning and uncertainty estimation.
@@ 7 Limitations & Future Work
The primary limitation of this study was its reliance on the normality assumption, which may not fully capture the potential of SEMF across diverse data distributions. Although in Appendix F we demonstrate that the framework can learn non-normal patterns, further investigation and exploration of SEMF under other distributions, such as uniform and log-normal, are needed. The computational complexity of the approach presents another significant challenge, as the current implementation can be optimized for large-scale applications. Additionally, while the CWR metric is useful, it implicitly assumes that a 1% drop in PICP equates to a 1% reduction in NMPIW, thus assuming a uniform distribution. Evaluating CWR under various distributional assumptions would provide a more comprehensive assessment of its implications. Finally, applying the same hyper-parameters across all datasets without specific tuning for incomplete data likely contributes to the observed decline in accuracy and robustness.
@@ Appendix C Optimal set of hyper-parameters
MultiXGBs and MultiMLPs benefit from early stopping to reduce computation time in complete and incomplete cases. Similarly, the baseline models for these instances use the same hyper-parameters for early stopping. Further, the number of epochs in the case of MultiMLPs is set as 1000, except for energy_efficiency and QSAR_fish_toxicity, where this is changed to 5000. Any model-specific hyperparameter we did not specify in this paper remains at the implementation’s default value (e.g., the number of leaves in XGBoost [50]). Along with the supplementary code, we provide two additional CSV files: one for the optimal hyperparameters of SEMF models and the other for the results and hyperparameters of all 330 runs.
@@ D.2 Impact of relative metrics for modeling
Appendix E Full results
@@
space_ga -1% 6% -7% 0.88 1.85 0.20 -10% -13% 0.60
space_ga -4% 5% -9% 0.89 1.85 0.20 -8% -10% 0.59
space_ga 1% 9% -8% 0.90 1.76 0.19 -10% -10% 0.62
space_ga -7% 5% -12% 0.90 1.92 0.21 -8% -10% 0.59
space_ga 8% 3% 4% 0.87 1.61 0.17 -9% -8% 0.60
cpu_activity 27% 7% 16% 0.88 0.37 0.07 21% 0% 0.98
cpu_activity 33% 11% 17% 0.89 0.37 0.07 23% -2% 0.98
cpu_activity 30% 10% 15% 0.88 0.36 0.07 20% -5% 0.98
cpu_activity 16% 9% 7% 0.89 0.38 0.07 20% -3% 0.98
cpu_activity 18% 9% 8% 0.90 0.38 0.07 22% 2% 0.98
naval_propulsion_plant 176% 8% 61% 0.96 0.39 0.12 -20% -14% 0.99
naval_propulsion_plant 151% 7% 58% 0.96 0.40 0.12 -9% -15% 0.99
naval_propulsion_plant 136% 8% 54% 0.96 0.40 0.12 -15% -12% 0.99
naval_propulsion_plant 156% 7% 58% 0.95 0.38 0.11 -12% -9% 0.99
naval_propulsion_plant 161% 9% 58% 0.96 0.38 0.11 -15% -11% 0.99
miami_housing 50% -5% 37% 0.82 0.46 0.06 3% -1% 0.91
miami_housing 65% -7% 44% 0.83 0.45 0.06 4% 1% 0.91
miami_housing 62% -10% 44% 0.81 0.44 0.06 7% 7% 0.90
miami_housing 62% -7% 43% 0.82 0.44 0.05 9% 5% 0.91
miami_housing 65% -9% 45% 0.81 0.44 0.05 -4% 0% 0.90
kin8nm 12% 4% 8% 0.89 1.88 0.37 -17% -20% 0.64
kin8nm 11% 5% 6% 0.89 1.91 0.37 -21% -22% 0.63
kin8nm 11% 5% 6% 0.90 1.92 0.38 -19% -21% 0.64
kin8nm 5% 7% -3% 0.91 2.06 0.40 -20% -23% 0.62
kin8nm 8% 3% 5% 0.88 1.91 0.37 -21% -24% 0.62
concrete_compressive_strength 7% 34% -26% 0.91 1.26 0.26 -34% -52% 0.86
concrete_compressive_strength 13% 22% -7% 0.89 1.25 0.25 -29% -52% 0.84
concrete_compressive_strength 13% 18% -5% 0.91 1.20 0.24 -21% -34% 0.86
concrete_compressive_strength 20% 31% -9% 0.93 1.27 0.26 -16% -22% 0.86
concrete_compressive_strength 14% 26% -10% 0.90 1.25 0.25 -28% -39% 0.83
cars 35% 14% 16% 0.89 0.67 0.11 -5% -1% 0.95
cars 16% 10% 5% 0.85 0.67 0.11 -4% -4% 0.95
cars 13% 20% -6% 0.91 0.78 0.13 -2% 0% 0.95
cars 12% 16% -4% 0.88 0.68 0.11 2% 5% 0.96
cars 16% 18% -2% 0.91 0.71 0.12 -7% -4% 0.95
energy_efficiency 255% 24% 64% 0.88 0.14 0.04 -50% -29% 1.00
energy_efficiency 320% 19% 71% 0.91 0.14 0.04 -3% -13% 1.00
energy_efficiency 257% 17% 67% 0.90 0.16 0.05 -21% -31% 1.00
energy_efficiency 322% 16% 73% 0.85 0.12 0.04 17% 16% 1.00
energy_efficiency 78% 34% 26% 0.89 0.22 0.06 -17% -22% 1.00
california_housing 24% 1% 18% 0.88 1.19 0.28 -2% -3% 0.81
california_housing 28% 1% 21% 0.88 1.16 0.28 0% -1% 0.82
california_housing 33% 0% 25% 0.89 1.19 0.28 0% 0% 0.81
california_housing 30% 1% 22% 0.87 1.16 0.28 -1% -1% 0.82
california_housing 38% -2% 29% 0.87 1.15 0.27 -1% -2% 0.81
airfoil_self_noise 68% 6% 37% 0.83 0.80 0.17 -41% -43% 0.89
airfoil_self_noise 134% 5% 55% 0.81 0.59 0.13 -18% -18% 0.93
airfoil_self_noise 13% -3% 15% 0.72 1.10 0.23 -109% -138% 0.78
airfoil_self_noise 38% 20% 13% 0.91 1.12 0.24 -73% -76% 0.85
airfoil_self_noise 51% -6% 38% 0.76 0.84 0.18 -82% -93% 0.86
QSAR_fish_toxicity 11% 17% -5% 0.77 1.48 0.23 3% 3% 0.53
QSAR_fish_toxicity 11% 9% 3% 0.75 1.42 0.22 8% 7% 0.58
QSAR_fish_toxicity 4% 8% -5% 0.77 1.48 0.23 4% 3% 0.57
QSAR_fish_toxicity 10% 9% 1% 0.77 1.45 0.23 -2% -1% 0.54
QSAR_fish_toxicity 13% 15% -1% 0.74 1.48 0.23 0% -6% 0.52

### v3

@@ 5.2 Missing Data
The results of models with 50% missing data are in Table 3. Note that relative metrics for SEMF are compared with the best result from any baseline imputer on that metric, regardless of how the imputer performed on the other metrics. Overall, the results are worse and less consistent than with complete data. MultiXGBs maintained good performance on some datasets, such as naval_propulsion_plant and energy_efficiency, but declined on others, like cpu_activity. MultiETs continued to exhibit mixed results, with only naval_propulsion_plant offering marginally better performance over the baseline. Similarly, MultiMLPs worked well on only one dataset, namely concrete_compressive_strength, while the other datasets had increased model uncertainty.
Table 3: Test results for all models with missing data at 95% quantiles aggregated over five seeds. For each metric, the mean and standard deviation of the performance across the seeds are separated by $\pm$. Performance over the baseline is highlighted in bold.
@@
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
@@
space_ga 0%$\pm$3-2%$\pm$12%$\pm$30.94$\pm$0.01 0.32$\pm$0.01 -7%$\pm$4-9%$\pm$30.45$\pm$0.06
cpu_activity -14%$\pm$90%$\pm$2-20%$\pm$140.94$\pm$0.03 0.18$\pm$0.02 -1%$\pm$21-17%$\pm$130.83$\pm$0.07
naval_propulsion_plant 18%$\pm$23-3%$\pm$212%$\pm$180.93$\pm$0.04 0.59$\pm$0.17 -14%$\pm$20-29%$\pm$350.77$\pm$0.05
miami_housing -18%$\pm$13-1%$\pm$1-25%$\pm$220.94$\pm$0.01 0.20$\pm$0.02 -14%$\pm$11-18%$\pm$100.72$\pm$0.10
kin8nm -9%$\pm$31%$\pm$1-13%$\pm$50.96$\pm$0.01 0.63$\pm$0.03 -3%$\pm$3-4%$\pm$40.40$\pm$0.02
concrete_compressive_strength 1%$\pm$9-2%$\pm$21%$\pm$100.95$\pm$0.01 0.57$\pm$0.02 -16%$\pm$12-20%$\pm$140.58$\pm$0.08
cars 9%$\pm$27-7%$\pm$46%$\pm$230.90$\pm$0.04 0.34$\pm$0.02 -30%$\pm$22-35%$\pm$240.66$\pm$0.15
energy_efficiency 26%$\pm$24-4%$\pm$418%$\pm$180.95$\pm$0.04 0.27$\pm$0.05 -186%$\pm$265-153%$\pm$1880.91$\pm$0.08
california_housing -6%$\pm$6-1%$\pm$1-8%$\pm$80.95$\pm$0.02 0.59$\pm$0.03 -4%$\pm$4-5%$\pm$50.69$\pm$0.05
airfoil_self_noise -24%$\pm$6-1%$\pm$2-32%$\pm$120.95$\pm$0.03 0.76$\pm$0.07 -37%$\pm$31-43%$\pm$360.41$\pm$0.12
QSAR_fish_toxicity 0%$\pm$8-4%$\pm$50%$\pm$110.91$\pm$0.05 0.50$\pm$0.11 -4%$\pm$4-2%$\pm$30.36$\pm$0.04
@@
space_ga -11%$\pm$20%$\pm$2-14%$\pm$50.95$\pm$0.01 0.34$\pm$0.02 -12%$\pm$4-13%$\pm$50.40$\pm$0.06
cpu_activity -15%$\pm$8-3%$\pm$2-18%$\pm$130.94$\pm$0.02 0.20$\pm$0.02 -21%$\pm$18-44%$\pm$140.82$\pm$0.07
naval_propulsion_plant 4%$\pm$16-4%$\pm$34%$\pm$180.95$\pm$0.03 0.73$\pm$0.13 -37%$\pm$21-95%$\pm$480.71$\pm$0.06
miami_housing -33%$\pm$100%$\pm$3-55%$\pm$240.95$\pm$0.01 0.27$\pm$0.02 -25%$\pm$10-39%$\pm$140.68$\pm$0.12
kin8nm -14%$\pm$11%$\pm$1-17%$\pm$20.95$\pm$0.01 0.66$\pm$0.02 -11%$\pm$2-13%$\pm$30.32$\pm$0.03
concrete_compressive_strength -20%$\pm$3-2%$\pm$4-27%$\pm$90.95$\pm$0.02 0.61$\pm$0.02 -37%$\pm$18-51%$\pm$220.48$\pm$0.06
cars -32%$\pm$12-4%$\pm$4-53%$\pm$320.93$\pm$0.02 0.38$\pm$0.05 -38%$\pm$15-32%$\pm$140.62$\pm$0.16
energy_efficiency -30%$\pm$28-4%$\pm$2-82%$\pm$1200.96$\pm$0.03 0.32$\pm$0.22 -69%$\pm$60-75%$\pm$430.94$\pm$0.03
california_housing -6%$\pm$4-2%$\pm$1-4%$\pm$40.96$\pm$0.01 0.70$\pm$0.03 -14%$\pm$3-17%$\pm$30.61$\pm$0.05
airfoil_self_noise -31%$\pm$4-3%$\pm$1-43%$\pm$90.94$\pm$0.02 0.75$\pm$0.05 -62%$\pm$51-89%$\pm$660.35$\pm$0.10
QSAR_fish_toxicity -7%$\pm$7-3%$\pm$3-7%$\pm$60.92$\pm$0.04 0.50$\pm$0.08 -9%$\pm$9-10%$\pm$90.39$\pm$0.09
@@
space_ga -25%$\pm$11-3%$\pm$2-34%$\pm$250.94$\pm$0.01 0.38$\pm$0.06 -40%$\pm$23-34%$\pm$150.21$\pm$0.23
cpu_activity -40%$\pm$90%$\pm$2-73%$\pm$240.95$\pm$0.01 0.25$\pm$0.02 -15%$\pm$19-23%$\pm$170.75$\pm$0.12
naval_propulsion_plant -43%$\pm$11-1%$\pm$2-89%$\pm$440.95$\pm$0.02 1.33$\pm$0.19 -108%$\pm$66-183%$\pm$195-0.03$\pm$0.47
miami_housing -36%$\pm$7-1%$\pm$1-57%$\pm$200.94$\pm$0.02 0.22$\pm$0.02 -25%$\pm$8-28%$\pm$110.67$\pm$0.10
kin8nm -19%$\pm$30%$\pm$0-24%$\pm$50.96$\pm$0.01 0.70$\pm$0.02 -16%$\pm$8-15%$\pm$90.37$\pm$0.04
concrete_compressive_strength 5%$\pm$10-3%$\pm$25%$\pm$100.94$\pm$0.02 0.54$\pm$0.04 -12%$\pm$5-12%$\pm$60.54$\pm$0.05
cars -22%$\pm$17-1%$\pm$4-38%$\pm$300.95$\pm$0.03 0.35$\pm$0.06 -28%$\pm$12-26%$\pm$130.66$\pm$0.10
energy_efficiency -7%$\pm$11-4%$\pm$3-11%$\pm$140.95$\pm$0.03 0.31$\pm$0.08 1%$\pm$9-1%$\pm$160.95$\pm$0.03
california_housing -24%$\pm$50%$\pm$1-34%$\pm$100.96$\pm$0.01 0.68$\pm$0.02 -5%$\pm$2-3%$\pm$30.61$\pm$0.06
airfoil_self_noise -8%$\pm$6-4%$\pm$3-7%$\pm$80.93$\pm$0.04 0.61$\pm$0.07 -15%$\pm$8-9%$\pm$80.55$\pm$0.18
QSAR_fish_toxicity -11%$\pm$5-2%$\pm$3-14%$\pm$80.92$\pm$0.04 0.53$\pm$0.11 -13%$\pm$7-14%$\pm$80.34$\pm$0.06
@@ 5.3 Discussion
Our results indicate that SEMF, when combined with XGBoost and MLPs, performs strongly on datasets with complete data. Both models produce better prediction intervals than traditional quantile regression methods in the complete case. MultiMLPs also deliver good point predictions for some datasets despite the experimental design prioritizing interval estimation over point accuracy. This could indicate either the effectiveness of the chosen $\sigma_{k}$, which (indirectly) led to narrower prediction intervals or underfitting of the MLPs. The truth may lie somewhere in between; SEMF’s sampling operation, akin to cross-validation, partially helps combat overfitting, ensuring robust results despite variations in the data. In our preliminary experiments, we observed that increasing the depth of the baseline MLPs did not help and eventually led to overfitting. ETs exhibit more mixed performance, highly dependent on the dataset’s characteristics. One possible explanation behind the larger variations in predictive power compared to other models may be ETs’ reliance on randomized splits, which introduces significant variability in the prediction process. Consequently, ETs may not benefit from the iterative sampling and refinement of predictions inherent in SEMF.
The robustness of SEMF diminishes when applied to datasets with missing data. Despite not offering improvements over the baseline in the presence of missing data, we have presented these results to transparently illustrate our framework’s current capabilities of handling missing inputs. From a theoretical standpoint, we believe there is value in further investigating how a single loss function that leverages EM’s missing data capabilities can be effectively applied. Given that SEMF performs well with complete data, a practical alternative might involve using the best imputation method for the data at hand before applying SEMF, effectively treating the data as complete. We expect our ablation study to perform well, given the complete data results.
An important observation is the stability of SEMF across different random seeds, as indicated by consistent PICP and NMPIW metrics (full results in Appendix E), contrasting the significant variability observed in the baseline models. This suggests that SEMF offers a more reliable performance framework. Additionally, it is worth noting that our experiments did not precisely tune for conformalized prediction intervals but used the non-conformalized (raw) quantiles from SEMF. The non-conformalized intervals offer less coverage but produce better CWR and NMPIW than the baseline (Appendix F). Certain experimental choices—such as processing the columns separately, using the same hyperparameters for both complete and incomplete data, and fixing the number of latent nodes per input ($m_{k}$)—may have further constrained the models’ ability to tailor their performance to the varying complexities of different datasets.
@@ 6 Conclusion
This paper introduces the Supervised Expectation-Maximization Framework (SEMF), a novel model-agnostic approach for generating prediction intervals in datasets with missing values. SEMF draws from the EM algorithm for supervised learning to devise latent representations that produce better prediction intervals than quantile regression. Due to SEMF’s iterative simulation technique, training and inference can be done with complete and incomplete data. A comprehensive set of 330 experimental runs on 11 datasets with three different model types showed that SEMF, in the case of complete data, outperforms quantile regression, particularly on complete datasets and when using XGBoost, which intrinsically lacks latent representations. The results of the missing data are less positive and require further investigation. This research underscores SEMF’s potential in various application domains and opens new avenues for further exploration of supervised latent representation learning and uncertainty estimation.
@@ 7 Limitations & Future Work
The primary limitation of this study was its reliance on the normality assumption, which may not fully capture the potential of SEMF across diverse data distributions. Although in Appendix G we demonstrate that the framework can learn non-normal patterns, further investigation and exploration of SEMF under other distributions, such as uniform, log-normal, and generalized extreme value distributions, are needed. The computational complexity of the approach presents another significant challenge, as the current implementation can be optimized for large-scale applications. Additionally, while the CWR metric is useful, it implicitly assumes that a 1% drop in PICP equates to a 1% reduction in NMPIW, thus assuming a uniform distribution. Evaluating CWR under various distributional assumptions would provide a more comprehensive assessment of its implications. Additionally, SEMF has only been evaluated on MCAR data and does not address missing at random cases (MAR), which require further investigation for real-world applicability. Finally, applying the same hyper-parameters across all datasets without specific tuning for incomplete data likely contributes to the observed decline in accuracy and robustness.
@@ Appendix C Optimal set of hyper-parameters
MultiXGBs and MultiMLPs benefit from early stopping to reduce computation time in complete and incomplete cases. Similarly, the baseline models for these instances use the same hyper-parameters for early stopping. Further, the number of epochs in the case of MultiMLPs is set as 1000, except for energy_efficiency and QSAR_fish_toxicity, where this is changed to 5000. Any model-specific hyperparameter we did not specify in this paper remains at the implementation’s default value (e.g., the number of leaves in XGBoost from [50]). Along with the supplementary code, we provide three additional CSV files: one for the results and hyperparameters of all 330 runs and the other two for the optimal hyperparameters of SEMF models, both raw (directly from SEMF) and conformalized.
@@ D.2 Impact of relative metrics for modeling
Appendix E Full conformalized results
@@
space_ga 2% -2% 4% 0.95 2.40 0.26 -10% -13% 0.60
space_ga 10% -1% 10% 0.95 2.34 0.25 -10% -10% 0.62
space_ga 8% -2% 9% 0.93 2.24 0.24 -9% -8% 0.60
space_ga 6% -2% 7% 0.94 2.38 0.26 -8% -10% 0.59
space_ga 7% -1% 7% 0.95 2.55 0.28 -8% -10% 0.59
cpu_activity 7% 1% 5% 0.95 0.49 0.09 22% 2% 0.98
cpu_activity 20% 1% 17% 0.93 0.45 0.09 20% -5% 0.98
cpu_activity 20% 0% 17% 0.93 0.45 0.09 21% 0% 0.98
cpu_activity 21% 1% 16% 0.94 0.45 0.09 23% -2% 0.98
cpu_activity 10% 2% 7% 0.95 0.47 0.09 20% -3% 0.98
naval_propulsion_plant 163% -1% 62% 0.95 0.36 0.11 -15% -12% 0.99
naval_propulsion_plant 190% 0% 66% 0.95 0.37 0.11 -20% -14% 0.99
naval_propulsion_plant 151% 0% 60% 0.96 0.40 0.12 -12% -9% 0.99
naval_propulsion_plant 185% -1% 65% 0.94 0.35 0.10 -9% -15% 0.99
naval_propulsion_plant 169% -1% 63% 0.96 0.37 0.11 -15% -11% 0.99
miami_housing -2% 0% -2% 0.95 0.99 0.12 4% 1% 0.91
miami_housing -10% 0% -10% 0.95 1.03 0.13 7% 7% 0.90
miami_housing -6% 1% -8% 0.95 0.99 0.12 9% 5% 0.91
miami_housing -7% 0% -8% 0.95 0.99 0.12 3% -1% 0.91
miami_housing -10% 0% -12% 0.95 1.05 0.13 -4% 0% 0.90
kin8nm 6% 0% 6% 0.94 2.27 0.45 -21% -24% 0.62
kin8nm 4% 1% 3% 0.95 2.39 0.47 -19% -21% 0.64
kin8nm 9% 1% 8% 0.94 2.20 0.43 -17% -20% 0.64
kin8nm 9% 1% 7% 0.95 2.36 0.46 -21% -22% 0.63
kin8nm 3% 1% 2% 0.94 2.33 0.46 -20% -23% 0.62
concrete_compressive_strength 40% -3% 31% 0.93 1.42 0.29 -16% -22% 0.86
concrete_compressive_strength 47% -6% 36% 0.91 1.48 0.30 -29% -52% 0.84
concrete_compressive_strength 59% -2% 39% 0.96 1.60 0.32 -21% -34% 0.86
concrete_compressive_strength 35% -3% 28% 0.95 1.48 0.30 -28% -39% 0.83
concrete_compressive_strength 24% 0% 19% 0.97 1.59 0.32 -34% -52% 0.86
cars 56% -4% 38% 0.91 0.70 0.12 -7% -4% 0.95
cars 36% -5% 30% 0.90 0.86 0.14 -4% -4% 0.95
cars 30% -3% 25% 0.91 0.76 0.13 2% 5% 0.96
cars 65% -1% 40% 0.93 0.75 0.13 -5% -1% 0.95
cars 16% -4% 17% 0.91 0.87 0.15 -2% 0% 0.95
energy_efficiency 165% -5% 64% 0.92 0.20 0.06 -50% -29% 1.00
energy_efficiency 288% 0% 74% 0.91 0.14 0.04 -3% -13% 1.00
energy_efficiency 217% -1% 69% 0.96 0.18 0.05 -21% -31% 1.00
energy_efficiency 253% -3% 73% 0.92 0.16 0.04 17% 16% 1.00
energy_efficiency 185% -10% 68% 0.91 0.23 0.07 -17% -22% 1.00
california_housing -5% 0% -6% 0.95 1.82 0.43 -2% -3% 0.81
california_housing 4% 0% 4% 0.95 1.72 0.41 -1% -2% 0.81
california_housing -4% 1% -5% 0.95 1.76 0.42 0% -1% 0.82
california_housing 1% 0% 1% 0.95 1.74 0.41 0% 0% 0.81
california_housing -3% 1% -4% 0.95 1.79 0.42 -1% -1% 0.82
airfoil_self_noise 15% -4% 16% 0.95 1.73 0.37 -82% -93% 0.86
airfoil_self_noise 45% 1% 30% 0.97 1.42 0.30 -18% -18% 0.93
airfoil_self_noise 5% 1% 4% 0.98 1.86 0.40 -73% -76% 0.85
airfoil_self_noise -6% 0% -7% 0.98 2.15 0.46 -109% -138% 0.78
airfoil_self_noise 46% -2% 33% 0.95 1.37 0.29 -41% -43% 0.89
QSAR_fish_toxicity 21% -2% 19% 0.88 2.12 0.33 3% 3% 0.53
QSAR_fish_toxicity 39% -9% 35% 0.85 1.97 0.31 8% 7% 0.58
QSAR_fish_toxicity 43% -3% 32% 0.89 2.21 0.34 4% 3% 0.57
QSAR_fish_toxicity 38% -3% 30% 0.89 2.11 0.33 -2% -1% 0.54
QSAR_fish_toxicity 29% -6% 28% 0.85 2.11 0.33 0% -6% 0.52